# Accurate and Diverse LLM Mathematical Reasoning via Automated PRM-Guided GFlowNets

### v2

@@ Abstract
Achieving both accuracy and diverse reasoning remains challenging for Large Language Models (LLMs) in complex domains like mathematics. A key bottleneck is evaluating intermediate reasoning steps to guide generation without costly human annotations. To address this, we first introduce a novel Process Reward Model (PRM) trained automatically using Monte Carlo Tree Search coupled with a similarity-based data augmentation technique, effectively capturing step-level reasoning quality. Leveraging this PRM, we then adapt Generative Flow Networks (GFlowNets) to operate at the reasoning step level. Unlike traditional reinforcement learning focused on maximizing a single reward, GFlowNets naturally sample diverse, high-quality solutions proportional to their rewards, as measured by our PRM. Empirical evaluation shows strong improvements in both accuracy and solution diversity on challenging mathematical benchmarks (e.g., +2.59% absolute accuracy on MATH Level 5 for Llama3.2-3B), with effective generalization to unseen datasets (+9.4% absolute on SAT MATH). Our work demonstrates the potential of PRM-guided, step-level GFlowNets for developing more robust and versatile mathematical reasoning in LLMs.
[…]
360
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable progress in various natural language tasks [4, 8], yet achieving robust and reliable mathematical reasoning remains a significant challenge [18, 11]. While LLMs have shown increasing proficiency in solving mathematical problems [6, 46], current approaches often prioritize accuracy on benchmark datasets [11, 41], potentially overlooking other crucial aspects of intelligent reasoning, such as the ability to explore and generate diverse solution strategies [25]. For LLMs to truly excel in mathematical domains and move beyond pattern recognition towards genuine understanding, they must not only arrive at correct answers but also exhibit the capacity to reason through problems in multiple, varied, and insightful ways [45, 38].
Traditional reinforcement learning methods like Proximal Policy Optimization [PPO; 33] have shown promise in improving LLM mathematical reasoning [44]. However, these methods inherently aim to maximize a single reward signal, often leading to the exploitation of a narrow set of solution strategies [51, 25]. This limitation becomes particularly critical when considering the development of robust and generally applicable problem-solving AI systems, where adaptability to novel situations and the exploration of diverse solution spaces are paramount.
Addressing the need for diverse reasoning requires effective guidance at the level of intermediate steps. However, obtaining reliable reward signals for these steps typically involves expensive human annotation. We overcome this limitation by first developing an automatically trained Process Reward Model [PRM; 38]. Our approach uniquely employs Monte Carlo Tree Search [MCTS; 20] combined with a novel similarity-based data augmentation technique to generate high-quality step-level reward signals without manual labeling. Given this automated and nuanced step-level reward mechanism provided by the PRM, we then propose leveraging Generative Flow Networks [GFlowNets; 1, 2] for fine-tuning. Unlike traditional reinforcement learning methods that often collapse to a single strategy, GFlowNets are designed to sample proportionally to rewards, naturally fostering solution diversity. Our key adaptation is operating GFlowNets at the reasoning step level rather than the token level – each state represents a partial solution, and actions generate complete reasoning steps – allowing the PRM reward to guide the exploration of diverse, high-quality reasoning paths effectively. Prior work using GFlowNets for LLM fine-tuning [13, 36] often utilized variants of the Subtrajectory Balance (SubTB) loss [21], which we adapt for our step-level framework.
@@
  • An efficient and scalable training methodology for mathematical reasoning, featuring an automatically trained Process Reward Model [PRM; 38] guided by Monte Carlo Tree Search [MCTS; 20] and a novel similarity-based data augmentation technique, effectively eliminating the need for expensive human annotations.
  • A step-level GFlowNet framework for mathematical reasoning, where states and actions correspond to meaningful solution components rather than individual tokens, enabling fine-grained control over reasoning path generation.
  • Comprehensive empirical validation showing both improved accuracy and enhanced solution diversity. Our approach achieves a 17.05% accuracy on MATH Level 5 using Llama3.2-3B (compared to 14.46% baseline and 15.32% with PPO), while generating measurably more diverse solution strategies. Additionally, our method generalizes to datasets unseen during the fine-tuning phase. For instance, we observe a relative increase of $14.32\%$ is observed on the SAT MATH dataset. Similar improvements are observed with an 8B model.
@@ 2 Background and Related Work
Mathematical Reasoning in LLMs has seen significant progress through various approaches, including chain-of-thought prompting [42], self-consistency [41], and reward modeling [19, 48]. While these methods have improved accuracy, they often lack mechanisms for promoting solution diversity. Recent work by Wang et al. [40] highlights the importance of diverse reasoning paths but focuses primarily on accuracy rather than explicitly encouraging diversity during training.
Chain-of-thought (CoT) approaches [42] and their variants such as Tree-of-Thought [43] and Graph-of-Thought [3] have demonstrated substantial improvements by encouraging models to articulate intermediate steps in their problem-solving process. However, the precise mechanisms underlying these improvements remain an active area of investigation, with some researchers questioning whether the benefits derive specifically from human-like task decomposition or simply from the additional computation afforded by generating more tokens [29]. Several enhancements to CoT approaches have been proposed, including using datasets of preference pairs of reasoning traces to finetune the CoT-generating model [17].
@@
Generative Flow Networks (GFlowNets) represent a novel framework for learning to sample from a desired distribution, offering advantages over traditional reinforcement learning approaches [1, 2]. GFlowNets operate on a directed acyclic graph (DAG) structure, where states $\mathcal{S}$ represent partial constructions and actions $\mathcal{A}$ represent transitions between states. This graph contains a unique source state $s_{0}$ with no parents and a sink state $s_{f}$ with no children. States that connect directly to $s_{f}$ are termed terminal states $\mathcal{X}$, each associated with a positive reward $R(x)>0$ for $x\in\mathcal{X}$.

### v3

@@ Abstract
Achieving both accuracy and diverse reasoning remains challenging for Large Language Models (LLMs) in complex domains like mathematics. A key bottleneck is evaluating intermediate reasoning steps to guide generation without costly human annotations. To address this, we first introduce a novel Process Reward Model (PRM) trained automatically using Monte Carlo Tree Search coupled with a similarity-based data augmentation technique, effectively capturing step-level reasoning quality. Leveraging this PRM, we then adapt Generative Flow Networks (GFlowNets) to operate at the reasoning step level. Unlike traditional reinforcement learning focused on maximizing a single reward, GFlowNets naturally sample diverse, high-quality solutions proportional to their rewards, as measured by our PRM. Empirical evaluation shows strong improvements in both accuracy and solution diversity on challenging mathematical benchmarks (e.g., +2.59% absolute accuracy on MATH Level 5 for Llama3.2-3B), with effective generalization to unseen datasets (+9.4% absolute on SAT MATH). Furthermore, we benchmark our PRM against existing open-source reward models, demonstrating superior alignment with reasoning quality and more consistent guidance for downstream generation. Our work demonstrates the potential of PRM-guided, step-level GFlowNets for developing more robust and versatile mathematical reasoning in LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable progress in various natural language tasks (Brown et al., 2020; Dubey et al., 2024), yet achieving robust and reliable mathematical reasoning remains a significant challenge (Lewkowycz et al., 2022; Hendrycks et al., 2021). While LLMs have shown increasing proficiency in solving mathematical problems (Cobbe et al., 2021; Yuan et al., 2023), current approaches often prioritize accuracy on benchmark datasets (Hendrycks et al., 2021; Wang et al., 2022), potentially overlooking other crucial aspects of intelligent reasoning, such as the ability to explore and generate diverse solution strategies (Naik et al., 2023). For LLMs to truly excel in mathematical domains and move beyond pattern recognition towards genuine understanding, they must not only arrive at correct answers but also exhibit the capacity to reason through problems in multiple, varied, and insightful ways (Yu et al., 2024; Uesato et al., 2022).
Traditional reinforcement learning methods like Proximal Policy Optimization (PPO; Schulman et al., 2017) have shown promise in improving LLM mathematical reasoning (Yao et al., 2023). However, these methods inherently aim to maximize a single reward signal, often leading to the exploitation of a narrow set of solution strategies (Ziegler et al., 2019; Naik et al., 2023). This limitation becomes particularly critical when considering the development of robust and generally applicable problem-solving AI systems, where adaptability to novel situations and the exploration of diverse solution spaces are paramount.
Addressing the need for diverse reasoning requires effective guidance at the level of intermediate steps. However, obtaining reliable reward signals for these steps typically involves expensive human annotation. We overcome this limitation by first developing an automatically trained Process Reward Model (PRM; Uesato et al., 2022). Our approach uniquely employs Monte Carlo Tree Search (MCTS; Luo et al., 2024) combined with a novel similarity-based data augmentation technique to generate high-quality step-level reward signals without manual labeling. Given this automated and nuanced step-level reward mechanism provided by the PRM, we then propose leveraging Generative Flow Networks (GFlowNets; Bengio et al., 2021; 2023) for fine-tuning. Unlike traditional reinforcement learning methods that often collapse to a single strategy, GFlowNets are designed to sample proportionally to rewards, naturally fostering solution diversity. Our key adaptation is operating GFlowNets at the reasoning step level rather than the token level – each state represents a partial solution, and actions generate complete reasoning steps – allowing the PRM reward to guide the exploration of diverse, high-quality reasoning paths effectively. Prior work using GFlowNets for LLM fine-tuning (Hu et al., 2023; Takase et al., 2024) often utilized variants of the Subtrajectory Balance (SubTB) loss (Madan et al., 2023), which we adapt for our step-level framework.
@@
  • Automated Process Reward Model: An efficient training methodology featuring an automatically trained PRM using Monte Carlo Tree Search and novel similarity-based data augmentation, eliminating the need for expensive human step-level annotations while achieving superior data efficiency through rollout reuse.
  • Step-Level GFlowNet Framework: An adaptation of GFlowNets to operate at complete reasoning steps rather than individual tokens, providing semantic coherence and enabling fine-grained quality control through step-wise PRM evaluation. This addresses key limitations of existing token-level approaches.
  • Strong Empirical Validation: Demonstrated improvements in both accuracy and solution diversity across challenging benchmarks, with particularly impressive generalization (+9.4% absolute on SAT MATH for 3B model), indicating that step-level diversity training captures transferable reasoning patterns.
@@ 2 Background and Related Work
Mathematical Reasoning in LLMs has seen significant progress through various approaches, including chain-of-thought prompting (Wei et al., 2022), self-consistency (Wang et al., 2022), and reward modeling (Lightman et al., 2023; Zhang et al., 2024). Process reward modeling, pioneered by Lightman et al. (2023), has evolved to include automated approaches: Zhang et al. (2024) introduced ReST-MCTS* for self-training via process reward guided tree search, while Guan et al. (2025) demonstrated that small language models can achieve state-of-the-art mathematical reasoning using MCTS-guided process rewards.
While these methods have improved accuracy, they often lack mechanisms for promoting solution diversity. Recent work by Wang et al. (2024b) highlights the importance of diverse reasoning paths but focuses primarily on accuracy rather than explicitly encouraging diversity during training.
Chain-of-thought (CoT) approaches (Wei et al., 2022) and their variants such as Tree-of-Thought (Yao et al., 2024) and Graph-of-Thought (Besta et al., 2024) have demonstrated substantial improvements by encouraging models to articulate intermediate steps in their problem-solving process. However, the precise mechanisms underlying these improvements remain an active area of investigation, with some researchers questioning whether the benefits derive specifically from human-like task decomposition or simply from the additional computation afforded by generating more tokens (Pfau et al., 2024). Several enhancements to CoT approaches have been proposed, including using datasets of preference pairs of reasoning traces to finetune the CoT-generating model (Lahlou et al., 2024).
@@
Generative Flow Networks (GFlowNets) represent a novel framework for learning to sample from a desired distribution, offering advantages over traditional reinforcement learning approaches (Bengio et al., 2021; 2023). GFlowNets operate on a directed acyclic graph (DAG) structure, where states $\mathcal{S}$ represent partial constructions and actions $\mathcal{A}$ represent transitions between states. This graph contains a unique source state $s_{0}$ with no parents and a sink state $s_{f}$ with no children. States that connect directly to $s_{f}$ are termed terminal states $\mathcal{X}$, each associated with a positive reward $R(x)>0$ for $x\in\mathcal{X}$.